• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name dalton

    Log entry time 13:07:27 on August 22, 2009

    Entry number 285194

    This entry is a followup to: 285179


    keyword=Compton slow ramp not working

    This is reminiscent of the last epics reboot that was done.

    At that time (Mon 10 Aug) we managed to fix all the issues except the
    temperature slow ramp. I believe that I tried the obvious step of
    pressing the button to turn on the slow ramp but I can't be sure.
    Anyway, when I later spoke to Sue Witherspoon this button was off and
    turning it on made the slow ramp work.

    Today we are unable to make the slow ramp work after trying many things
    including: power cycling the laser, using the ramp button, and altering
    the "Tracking Level (V)". It is also possible now to change the
    temperature of the laser while it is locked, using the increment and
    decrement button - something which I think didn't used to be possible
    before (and must be related.) I wonder if this is all related to some
    default configuration that is loaded after a reboot (which later got
    changed after the previous reboot allowing us to turn the ramp on.)

    I will speak to Sue on Monday to try and make this slow ramping
    functional again.

    A copy of this log entry has been emailed to: nanda@jlab.org, camsonne@jlab.org